A bank teller has P20 and P50 bills. The number of P50 bills is 7 less than twice the number of P20 bills. If she has a total of P1090. How many P20 and P50 bills are there?
Answer by macston(5194)

.
X=number of P20 bills; Y=number of P50 bills=2X-7
.
P20(X)+P50(Y)=P1090
20x+50(2X-7)=1090
20X+100X-350=1090
120X=1440
X=12
ANSWER 1: There are twelve P20 bills.
.
Y=2X-7=2(12)-7=24-7=17
ANSWER 2: There are seventeen P50 bills.
.
CHECK:
P20(X)+P50(Y)=P1090
P20(12)+P50(17)=P1090
P240+P850=P1090
P1090=P1090